OpenStudy (anonymous):

he point (3, 4) is on the terminal side of an angle in standard position. Determine the exact value of the cosine function of this angle.

OpenStudy (anonymous):

you have a nice 3 - 4 - 5 right triangle to work with 

OpenStudy (anonymous):

so you know all the necessary ratios. \[\sin(\theta)=\frac{4}{5},\cos(\theta)=\frac{3}{5} \] etc

OpenStudy (anonymous):

oh does cos = adj/opp

OpenStudy (anonymous):

for a right triangle cosine is adjacent over hypotenuse